OpenStudy (anonymous):

I will meadel and fan!!!!!!!!!!!!!!!!! round the mixednumber to the nearest 1/2. 4 2/3 A.4 B. 4 1/2 C. 5

OpenStudy (igreen):

4 2/3 is approximately 4.67. Is 4.67 closer to 4.5 or 5?

OpenStudy (anonymous):

5

OpenStudy (anonymous):

ARE U THERE

OpenStudy (igreen):

\(4.67 - 4.5 = 0.17\) \(5 - 4.67 = 0.33\) Which is smaller \(0.17\) or \(0.33\)?

OpenStudy (anonymous):

0.17

OpenStudy (igreen):

Right, so \(4.5\) is closer to \(4.67\).

OpenStudy (anonymous):

OKA

OpenStudy (anonymous):

SO WHAT IS MY ANSWER

OpenStudy (igreen):

\(4 \dfrac{1}{2}\) is the same thing as \(4.5\).

OpenStudy (igreen):

So \(4 \dfrac{1}{2}\) is your answer..

OpenStudy (anonymous):

O YEA SORRY I THOUGHT THAT WAS THE ANSWER

OpenStudy (anonymous):

I JUST WASENT SURE

OpenStudy (igreen):

Okay